# Regular AdS3 black holes from regularized Gauss-Bonnet coupling

### v1

@@ 3 Solution in [14] revisited
    with which the norm of the Weyl vector is $W^{2}=N-1$. Setting $N=1$, which is forced by the consistency of the equations, one directly studies the field configurations where the Weyl vector is null.
@@
    Equations of motion are now decoupled. Solve $w_{0}$ first. Then, the equation for the metric function can be easily found.
@@ 4 Regular Black Hole Solutions
$L_{\text{red}}=L^{\text{GR}}_{\text{red}}+\ell^{2}\left(L_{\text{red}}^{[B]}-\kappa L_{\text{red}}^{[C]}\right).$ (4.4)
@@
By the method of [44] that we have outlined in the previous section, studying the solutions of the undeformed ($\kappa=1$) and the deformed ($\kappa\neq 1$) theory is a straightforward exercise. We now present regular black hole solutions and their charged versions by coupling the theories to Born–Infeld electrodynamics.
4.1 Undeformed theory ($\mathbf{\kappa=1}$)
@@
$b_{0}=\frac{q_{b}}{r}-\frac{f}{2r}\quad\quad c_{0}=\frac{q_{c}}{r}-\frac{f}{2r},$ (4.6)
@@
and the only other independent curvature invariants in 3d ($R^{\mu}_{\ \nu}R^{\nu}_{\ \mu}$ and $R^{\mu}_{\ \rho}R^{\rho}_{\ \nu}R^{\nu}_{\ \mu}$) [45, 46, 47] are also finite at the origin. Therefore, we conclude that it is a regular black hole solution with two primary hairs. Note that for $q_{b}=q_{c}$, one obtains a stealth BTZ black hole solution.
4.2 Undeformed theory ($\mathbf{\kappa=1}$) coupled to Born–Infeld electrodynamics
We now couple the theory to Born–Infeld electrodynamics with the Lagrangian
$\mathcal{L}_{\text{BI}}=b^{2}\left(1-\sqrt{1+\frac{F_{\mu\nu}F^{\mu\nu}}{2b^{2}}}\,\right)$ (4.10)
where $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the field strength tensor of the $U(1)$ gauge field $A_{\mu}$. In the $b\to\infty$ limit of the Born-Infeld parameter, it reduces to the Lagrangian of Maxwell’s theory $\mathcal{L}_{\text{M}}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$. By taking
@@ 4.2 Undeformed theory ($\mathbf{\kappa=1}$) coupled to Born–Infeld electrodynamics
the scalar potential can be found as
@@
$R=\frac{3\left(2q_{b}^{2}+m\sqrt{\kappa}\left|(\kappa-1)\ell^{2}q_{b}\right|\right)}{2(\kappa-1)\ell^{2}q_{b}^{2}}+\mathcal{O}(r^{2}),$ (4.15)
@@
4.3 Deformed theory ($\mathbf{\kappa\neq 1}$)
@@
4.4 Deformed theory ($\mathbf{\kappa\neq 1}$) coupled to Born–Infeld electrodynamics
@@
$f$ $=-\frac{r^{2}}{2(\kappa-1)\ell^{2}}\times$ (4.21)
@@ 5 Conclusion
With these solutions, it is possible to ask questions about the black hole interior for a regular black hole in 3d. Previous studies show that regular AdS3 black holes are geodesically complete and tidal forces experienced by observers are finite everywhere [13, 38]. The same behavior is expected for the solutions presented here. Because of the hair parameters, the thermodynamical properties of the black hole solutions given here might be particularly interesting.

### v2

@@ 3 Solution in [14] revisited
    with which the norm of the Weyl vector vanishes ($W^{2}=0$). Thanks to this choice, one can easily get the solution where there is no additional integration constant (other than the hair parameter) appear in the vector components. This is equivalent to setting the additional constant to zero at the end by using the global symmetry of the reduced action, as explained before.
@@
    Then, one sees that the consistency of equations of motion implies $N=1$ and they now become decoupled. Solving $w_{0}$ first, the metric function $f$ can be easily found.
@@ 4 Regular black hole solutions
$L_{\text{red}}=L^{\text{EH}}_{\text{red}}+\ell^{2}\left(L_{\text{red}}^{[B]}-\kappa L_{\text{red}}^{[C]}\right).$ (4.4)
@@
By the method of [44] that we have outlined in the previous section, studying the solutions of the undeformed ($\kappa=1$) and the deformed ($\kappa\neq 1$) theory is a straightforward exercise. We now present regular black hole solutions and their charged versions by coupling the theories to Born–Infeld electrodynamics. In Appendix A, charged black hole solutions obtained by coupling to Maxwell’s electrodynamics are given. Due to the logarithmic form of the Coulomb potential in 3d, they possess a curvature singularity.
4.1 Undeformed theory ($\kappa=1$)
@@
$b_{0}=\frac{q_{b}}{r}-\frac{f}{2r},\quad\quad c_{0}=\frac{q_{c}}{r}-\frac{f}{2r},$ (4.6)
@@
and the only other independent curvature invariants in 3d ($R^{\mu}_{\ \nu}R^{\nu}_{\ \mu}$ and $R^{\mu}_{\ \rho}R^{\rho}_{\ \nu}R^{\nu}_{\ \mu}$) [45, 46, 47] are also finite at the origin. Therefore, we conclude that it is a regular black hole solution with two primary hairs. Note that $q_{b}=q_{c}$ makes $B=C$ on-shell, and one obtains a stealth BTZ black hole solution.
4.2 Undeformed theory ($\kappa=1$) coupled to Born–Infeld electrodynamics
We now couple the theory to Born–Infeld electrodynamics with the Lagrangian [48]
$\mathcal{L}_{\text{BI}}=b^{2}\left(1-\sqrt{1+\frac{F_{\mu\nu}F^{\mu\nu}}{2b^{2}}}\,\right),$ (4.10)
where $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the field strength tensor of the $U(1)$ gauge field $A_{\mu}$. In the $b\to\infty$ limit of the Born-Infeld parameter, it reduces to the Lagrangian of Maxwell’s theory in (A.1). By assuming
@@ 4.2 Undeformed theory ($\kappa=1$) coupled to Born–Infeld electrodynamics
the scalar potential solving the equation of motion can be taken as [49]
@@
$R=\frac{3\left(2q_{b}^{2}+m\sqrt{\kappa}\,\ell^{2}\left|(\kappa-1)q_{b}\right|\right)}{2(\kappa-1)\ell^{2}q_{b}^{2}}+\mathcal{O}(r^{2}),$ (4.15)
@@
4.3 Deformed theory ($\kappa\neq 1$)
@@
4.4 Deformed theory ($\kappa\neq 1$) coupled to Born–Infeld electrodynamics
@@
$f$ $=-\frac{r^{2}}{2(\kappa-1)\ell^{2}}\times$
@@ 5 Conclusion
With these solutions, it is possible to ask questions about the black hole interior for a regular black hole in 3d. Previous studies show that regular AdS3 black holes are geodesically complete and tidal forces experienced by observers are finite everywhere [13, 38]. The same behavior is expected for the solutions presented here. Because of the hair parameters, the thermodynamical properties of the black hole solutions given here might be particularly interesting (see [50, 51] for previous 3d examples).
@@
Appendix A Coupling to Maxwell’s electrodynamics
In this appendix, we give the resulting charged black hole solutions when the undeformed ($\kappa=1$) and the deformed ($\kappa\neq 1$) theory are coupled to Maxwell’s electrodynamics with the Lagrangian
$\mathcal{L}_{\text{M}}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}.$ (A.1)
Taking the vector field as in (4.11), one finds the Coulomb potential
$\phi=-q_{e}\log(r).$ (A.2)
The components of the Weyl vectors are given in (4.5, 4.6), and the metric functions are as follows
$\kappa=1:\quad\quad f$ $=\frac{r^{2}\left[-\Lambda_{0}r^{2}-m-\frac{q_{e}^{2}}{2}\log(r)\right]+4\ell^{2}(q_{b}^{2}-q_{c}^{2})}{r^{2}+4\ell^{2}(q_{b}-q_{c})},$ (A.3)
$\kappa\neq 1:\quad\quad f$ $=-\frac{r^{2}}{2\ell^{2}(\kappa-1)}\times$
$\left(1-\sqrt{1-4\Lambda_{0}(\kappa-1)\ell^{2}+\frac{4(\kappa-1)^{2}\ell^{4}(4q_{b}^{2}-\kappa mr^{2})}{\kappa r^{4}}-\frac{2(\kappa-1)q_{e}^{2}\ell^{2}\log(r)}{r^{2}}}\,\right).$ (A.4)
From the $r\to 0$ limit, the behavior of the Ricci scalar is found as
$\kappa=1:\quad\quad R$ $=\frac{12(q_{b}+q_{c}+m)+5q_{e}^{2}+6q_{e}^{2}\log(r)}{8\ell^{2}(q_{b}-q_{c})}+\mathcal{O}(r^{2}),$ (A.5)
$\kappa\neq 1:\quad\quad R$ $=\frac{12\kappa(\kappa-1)m\ell^{2}+5\kappa q_{e}^{2}+24\sqrt{\kappa}q_{b}+6\kappa q_{e}^{2}\log(r)}{8(\kappa-1)\sqrt{\kappa}\,\ell^{2}q_{b}}+\mathcal{O}(r^{2}).$ (A.6)
We see that these solutions are singular at the origin.